Computer Science 271 Software I

Dr. Stephen Bloch office 112 Alumnæ Hall phone 877-4483

                email sbloch@boethius.adelphi.edu

office hours: Monday, Wednesday, and 4:20-5:00 on Tuesday and Thursday. Other times by appointment.

September 12, 1997

1 Prerequisites

This course assumes you have taken and passed CSC 171 (“Introduction to Computer Programming”) and CSC 172 (“Introduction to Algorithms and Data Structures”), and that you either have taken or are now taking CSC 270 (“Survey of Programming Languages”). This course will demand a lot of time programming, testing, and debugging in C, particularly in the latter half of the semester. If you have not taken CSC 270 but are familiar with the C or C++ programming language, you should have no problem; otherwise, you should reconsider taking this course.

2 Subject Matter

This course will explain the basic concepts common to many computer operating systems and introduce you to a great number of tools and features provided by operating systems, especially Unix tm, which is currently the most popular at academic and research institutions and coming into widespread use elsewhere.

The Unix tm operating system has more personality than most, not because of how it works but because the people who built (and continue to build) it left such a distinctive mark on it with their senses of beauty, humor, and responsible behaviour. These have led to a system which, once you’ve grasped its fundamental philosophy, allows you both to make efficient use of your time and to have fun at it.

The creators of Unix tm used computers all day, every day, and wanted to use them as efficiently as possible. These people loved elegance and simplicity, and hated re-inventing the wheel, so they built a number of simple tools that could be combined as needed in billions of ways. They also loved whimsy, so you will find a lot of puns and jokes built into the system. On the other hand, they believed strongly in personal responsibility, so destructive pranks, invasions of privacy, and insults are strongly discouraged — usually not by rules, but by peer pressure. Unix tm has evolved from a computer system to a philosophy and finally to a community. I hope, among other things, to make you productive and civilized members of that community.

A disadvantage of building the whole system out of simple tools is that you need to learn a lot of simple tools before you can do anything. So in the first half of the semester you’ll read a lot of documentation and try your skill on small examples. In the second half of the semester the emphasis will shift from reading about Unix tm tools to actually using them.

Many of the people in this class are taking CSC 270 concurrently. You’ll need to learn the same Unix tm tools and commands for both courses, but whereas for 270 it will suffice to learn only their simplest, most basic uses, in this class we’ll discuss each tool and command in greater detail.

3 How to Learn Unix ^{tm}

This is not a theoretical course; it is a very practical, hands-on course. To learn the stuff I expect you to learn, you must spend a lot of time on the computer trying things. Every time you read about a new command, or hear me describe one in lecture, try it. Invent new ways to use it, beyond what I or the textbook have described. Play with it. The best way to learn Unix tm is to use it until you don’t even have to think about the most common commands — your fingers know them by heart.

4 Texts

The main text for this course is An Introduction to Unix, with X and the Internet, by Paul S. Wang. I expect you to read most of the book by the end of the semester, about twenty pages per class meeting. (We’ll skip a few sections, most notably the section on network programming using sockets — I’d love to cover this, and encourage you to read about it on your own, but there isn’t enough time in the semester.) I may also give some reading assignments in class handouts, on the World Wide Web, by email, in magazines, and in the CSC 270 textbook.

You are responsible for everything in the reading assignments, whether or not I discuss it in a lecture.

5 Grading

As I write this (a week before classes start), I envision four homework assignments, each worth 15% of the semester grade, a midterm worth 15%, and a final exam worth 25%. I may change this somewhat as the semester unfolds.

Exams must be taken at the scheduled time, unless arranged in advance or prevented by a documented medical or family emergency. If you have three or more exams scheduled on the same date, or a religious holiday that conflicts with an exam or assignment due date, please notify me in writing within the first two weeks of the semester in order to receive due consideration. Exams not taken without one of the above excuses will get a grade of 0.

Homework and programming assignments will be accepted late, with a penalty of 1/3 per 24 hours or portion thereof after they’re due. An hour late is 33% off, 25 hours late is 67% off, and after 48 hours don’t bother turning it in. It’s still a good idea to do as much of it as you can, however, because I’ll assume on the exams that you’ve done the homework.

Programs are not abstract works of art, they are supposed to run and solve real problems. So if I get a program that doesn’t compile or run, or a program that has little or nothing to do with the problem I assigned, I will give it a zero, no matter how much time you put into it. Don’t bother turning in a program you haven’t tested yourself.

6 Ethics

Most of the assignments in this class are to be done individually. You may discuss general approaches to a problem with classmates, but you may not copy large pieces of programs or homework solutions. If you do, all the students involved will be penalized.

All work on an exam must be entirely the work of the one person whose name is at the top of the page. If I have evidence that one student copied from another on an exam, both students will be penalized; see above.

7 Schedule

This class meets every Tuesday and Thursday from 12:15 to 1:30 PM, except on University holidays or if I cancel class. All dates in the following schedule are tentative, except those fixed by the University; if some topic listed here as taking one lecture in fact takes two lectures to cover adequately, or vice versa, the schedule will shift.

All reading assignments below are from the Wang textbook unless stated otherwise. I expect you to have read the specified chapters in the textbook before the lecture that deals with that topic; this way I can concentrate my time on answering questions and clarifying subtle or difficult points in the textbook, rather than on reading to you, which will bore both of us. Please read ahead!

Date(s) Assignment Reading Lecture Subject 2 Sept Accounts, passwords, what is Unix tm? 4 Sept HW1 Intro, 1–1.5 The Basics of Everything 9 Sept 1.6–1.14, info info More Basic Commands 11 Sept 2 The vi text editor 12 Sept Last day to add courses

16 Sept  3  Windows and keyboards and mice, oh my!
18 Sept  4  More on electronic mail
23 Sept  HW1 due 5–5.7  Commands and Filters
25 Sept  HW2  5.8–5.14  Filters and Pipes

26 Sept Last day to drop courses 30 Sept 6 The Shell Game 2 Oct Rosh Hashanah, no classes

7 Oct  7–7.7  Automating the Shell
9 Oct  7.8–7.18  More of the same
14 Oct  8  The Unix tm file system
16 Oct  9–9.7  Using the Net

21 Oct HW2 due 9.8–9.12, Web pages More Net stuff; review for midterm 23 Oct Midterm exam 28 Oct Discuss midterm 30 Oct HW3 10–10.5, man pages Shells, programs, and stdio 31 Oct Last day to withdraw from classes 4 Nov 10.6–10.9, info gdb Debugging C programs

6 Nov  11–11.6  Low-level Unix tm I/O
11 Nov  11.7–11.15  Process control
13 Nov  HW3 due 11.16–12.3  More on process control and I/O
18 Nov  HW4  handouts, man flex The flex scanner generator
20 Nov  13–13.7  The make utility

25 Nov make meets flex 27 Nov Thanksgiving — no classes 2 Dec man rcsintro, etc. The rcs version-control utility 4 Dec 14 L^{A}TEX and document preparation 9 Dec Catch up and review for final 11 Dec HW4 due Catch up and review for final 16 Dec 1:00–3:00 PM, Final Exam